A motor boat goes downstream twice as fast as upstream. Find the power speed of the boat if the water flows at 6km/h.
This is what I've came up with:
Let x be the time of going upstream
(and I'm terribly stuck) Answer by mananth(16949) 
 upstream speed be x
downstream speed =2x
boat speed = y
upstream
y-6=x............(1)
downstream
y+6=2x
y=2x-6
2x=y+6
x=(y+6)/2.................(2)
equate both
(y+6)/2 = y-6
y+6=2y-12
y=18
18km/h speed of boat